# Missing link between the 2D Quantum Hall problem and 1D quasicrystals

### v2

@@ II The Fibonacci, Quantum Hall and Fibonacci-Hall models
A 2D Quantum Hall Hamiltonian $H^{2D}$ is now defined. To reproduce the Fibonacci winding property, a fictitious dimensionless flux $\phi^{S}=Ba^{2}/\phi_{0}$ is introduced, where $B$ is a fictitious magnetic field perpendicular to the xyx-y plane, $a$ the side length and $\phi_{0}$ the flux quantum. The superscript “S” indicates that this flux arises due to the winding property induced by the strip $S$. The flux $\phi_{S}$ has a different rational value for each approximant, and is set by the winding of the strip in the 2D lattice. Fig.1d) is an equivalent compact representation of an periodic approximant corresponding to a slope $p/q$, showing the windings of the rational strip around the unit torus, $q$ times and $p$ along the $x$ and $y$ directions respectively. To fix $\phi_{S}$ we require a) that the total flux for the path shown in Fig.1c), namely, $(p+q)\phi^{S}$ be an integer. This means that $\phi^{S}\propto(p+q)^{-1}$. We require next b) that it lead to the hierarchical structure of Fibonacci chain spectra, as seen by explicit computation in a perturbative limit [18, 19, 20]. In the Quantum Hall problem, we recall a result concerning sub-bands of the spectrum due to Hofstadter [21]. It states that, for a rational flux $f$, the effective flux governing sub bands is $\phi^{\prime}=[f^{-1}]$. Now, setting $f=F_{k-1}/F_{k}$, then using the recursion relation for Fibonacci numbers, it is easy to see that
@@
To resume, $H^{FH}$ depends on several parameters. The approximant size $F_{k}$, firstly, fixes the value of flux $\phi^{S}$. The explicit parameters of the Hamiltonian are $t_{a}$, $t_{b}$ and $\epsilon$ of which only two are independent, (we can set $t_{a}=1$ without loss of generality). We will not discuss the extension of the model to include onsite potentials, but this could be done in an analogous way. There is an additional implicit parameter namely, the position of the selection strip. This parameter has no incidence on the band structure when periodic boundary conditions are applied but does play a role in the energies of the edge states in an open system. Nb. the $H^{FH}$ model for zero “flux” is topologically trivial but expected to have electronic properties different from the square Fibonacci model [23, 24] (see S.I.).
@@ III Phase diagram of the Fibonacci-Hall model
is well obeyed by the data (see S.I.). This Lifshitz transition is indicated in Fig.2b) by the dashed blue lines demarcating the topological and trivial regions.
@@ IV Edge states for $\delta\neq 0$ and Lifshitz transition
In open systems the bulk-edge correspondence leads to edge modes, which can be used as markers for the various topological transitions in the $\epsilon-\delta$ plane. We considered square unit cells (as discussed in the S.I.) and solved for the eigenstates of Eq.II when the long edges are left open, and periodic boundary conditions are applied along the $y^{\prime}$ direction. Here we study edge states in the main gap for fixed finite $\delta$. When $\epsilon=1$, one is in the topological sector, with an edge state. As $\epsilon$ is reduced, a topological transition occurs, with disappearance of the edge state and simultaneously a change of the Chern number.

### v3

@@ II The Fibonacci, Quantum Hall and Fibonacci-Hall models
A 2D Quantum Hall Hamiltonian $H^{2D}$ is now defined. To reproduce the Fibonacci winding property, a fictitious dimensionless flux $\phi^{S}=Ba^{2}/\phi_{0}$ is introduced, where $B$ is a fictitious magnetic field perpendicular to the xyx-y plane, $a$ the side length and $\phi_{0}$ the flux quantum. The superscript “S” indicates that this flux arises due to the winding property induced by the strip $S$. The flux $\phi_{S}$ has a different rational value for each approximant, and is set by the winding of the strip in the 2D lattice. Fig.1d) is an equivalent compact representation of an periodic approximant corresponding to a slope $p/q$, showing the windings of the rational strip around the unit torus, $q$ times and $p$ along the $x$ and $y$ directions respectively. To fix $\phi_{S}$ we require a) that the total flux for the path shown in Fig.1d), namely, $(p+q)\phi^{S}$ be an integer. This means that $\phi^{S}\propto(p+q)^{-1}$. We require next b) that it lead to the hierarchical structure of Fibonacci chain spectra, as seen by explicit computation in a perturbative limit [18, 19, 20]. In the Quantum Hall problem, we recall a result concerning sub-bands of the spectrum due to Hofstadter [21]. It states that, for a rational flux $f$, the effective flux governing sub bands is $\phi^{\prime}=[f^{-1}]$. Now, setting $f=F_{k-1}/F_{k}$, then using the recursion relation for Fibonacci numbers, it is easy to see that
@@
To resume, $H^{FH}$ depends on several parameters. The approximant size $F_{k}$, firstly, fixes the value of flux $\phi^{S}$. The explicit parameters of the Hamiltonian are $t_{a}$, $t_{b}$ and $\epsilon$ of which only two are independent, (we can set $t_{a}=1$ without loss of generality). We will not discuss the extension of the model to include onsite potentials, but this could be done in an analogous way. There is an additional implicit parameter namely, the position of the selection strip. This parameter has no incidence on the band structure when periodic boundary conditions are applied but does play a role in the energies of the edge states in an open system. Nb. the $H^{FH}$ model for zero “flux” is topologically trivial but expected to have electronic properties different from the square Fibonacci model [23, 24] (a compact expression for our Hamiltonian using the so-called conumbering scheme [25] is given in the S.I. [26] ).
@@ III Phase diagram of the Fibonacci-Hall model
is well obeyed by the data (see S.I. [26]). This Lifshitz transition is indicated in Fig.2b) by the dashed blue lines demarcating the topological and trivial regions.
@@ IV Edge states for $\delta\neq 0$ and Lifshitz transition
In open systems the bulk-edge correspondence leads to edge modes, which can be used as markers for the various topological transitions in the $\epsilon-\delta$ plane. We considered square unit cells (as discussed in the S.I. [26]) and solved for the eigenstates of Eq.II when the long edges are left open, and periodic boundary conditions are applied along the $y^{\prime}$ direction. Here we study edge states in the main gap for fixed finite $\delta$. When $\epsilon=1$, one is in the topological sector, with an edge state. As $\epsilon$ is reduced, a topological transition occurs, with disappearance of the edge state and simultaneously a change of the Chern number.